# Sampled-Data Primal-Dual Gradient Dynamics in Model Predictive Control

### v2

@@ 3.1 Assumptions
For $S^{\mathcal{P}}(x):=\frac{1}{2}x^{\top}x$, there exists $Q=Q^{\top},S,R=R^{\top}$ such that
@@
This assumption can be satisfied for any linear plant with matrices such as
$H^{\mathcal{P}}_{\rm c}$ $=\begin{bmatrix}(A_{\rm c}+A_{\rm c}^{\top})/2&B_{\rm d}/2\\ B_{\rm d}^{\top}/2&O\end{bmatrix}.$
To satisfy the stability condition in iMPC, however, $Q$ must be at least negative semi-definite. If it is not, it can be satisfied by pre-controlling the plant (1) with linear feedback *1*1*1Just apply $u=Kx+v$ with an appropriate gain $K$ and regard $v$ as a new input.. This allows us to apply our method to any controllable linear systems.
@@ 3.2 Instant MPC
holds for any $w,x$*2*2*2If $f(w)=\frac{1}{2}{\left\|{w}\right\|}^{2}_{P}\ (P=P^{\top},\sigma I\preceq P)$, then $\sigma I$ can be replaced as $P$ hereafter to reduce conservatism..
@@ 4.2 Assumptions
There exists matrix $H^{\mathcal{P}}_{\rm d}$ and positive semi-definite matrix $P^{\mathcal{P}}_{\rm d}$ such that
@@
Assumption 11 is an extension of Assumption 4, and for a linear plant $\mathcal{P}_{\rm d}$, there always exists matrices satisfying the assumption, e.g.,
$H^{\mathcal{P}}_{\rm d}$ $=\frac{1}{\Delta t}\begin{bmatrix}(A_{\rm d}+A_{\rm d}^{\top})/2-I&B_{\rm d}/2\\ B_{\rm d}^{\top}/2&O\end{bmatrix},$
$P^{\mathcal{P}}_{\rm d}$ $=\frac{1}{2(\Delta t)^{2}}\begin{bmatrix}A_{\rm d}^{\top}-I\\ B_{\rm d}^{\top}\end{bmatrix}\begin{bmatrix}A_{\rm d}-I&B_{\rm d}\end{bmatrix}.$
@@
In addition to Assumption 8, $f$ is $\rho$–smooth convex*3*3*3A function $f(x)$ is called $\rho$–smooth convex function if it satisfies ${\left\|{\nabla{f(x)}-\nabla{f(y)}}\right\|}\leq\rho{\left\|{x-y}\right\|}$ for all $x,y$. and $\mu_{0}\geq 0$.
@@ 4.3 Controller Design
If $z_{k}\neq 0$, small $\bar{\gamma}_{k}$ satisfies (a) because $-{\left\|{z_{k}}\right\|}^{2}_{-\bar{H}_{\rm d}}<0$. If $z_{k}=0$ and $\mu_{k}\neq 0$, $\Delta V_{k}\leq\zeta\Delta t{\left(a(\mu_{k},0)\gamma_{k}+b(\mu_{k},0)\right)}\gamma_{k}$ holds from (11) and small $\bar{\gamma}_{k}$ satisfies (a) because $b(\mu_{k},0)<0$. If $z_{k}=0$ and $\mu_{k}=0$, this means $(w_{k},\mu_{k},\lambda_{k},x_{k})=(0,0,0,0)$ and (a) is trivial because $\Delta V_{k}=0\ \forall\ \gamma_{k}\in(0,1]$. To summarize what is required to guarantee stability, we need to find $\delta>0$ that satisfies (10) and determine $\gamma_{k}$ satisfying (a) at each time step. The existence of $\delta$ can be confirmed by solving an LMI, e.g., $\min_{\{\delta,{\varepsilon}\}}{\varepsilon}\ \text{s.t. }{\varepsilon}I\succeq\bar{H}_{\rm d}$. We can see from (11) that $\gamma_{k}$ can be determined by $\gamma_{k}=\min(1,c\bar{\gamma}_{k}),\ c\in(0,1)$ with deriving $\bar{\gamma}_{k}$ as the positive solution of an equation*4*4*4An example of functions $a,b$ is shown in the proof of Lemma 17 (see Appendix). Replacing $-{\left\|{z_{k}}\right\|}^{2}_{-\bar{H}_{\rm d}}$ as ${\varepsilon}{\left\|{z_{k}}\right\|}^{2}$ using the LMI solution ${\varepsilon}$ is possible to avoid a large matrix computation. $a(\mu_{k},z_{k})\gamma_{k}^{2}+b(\mu_{k},z_{k})\gamma_{k}-{\left\|{z_{k}}\right\|}^{2}_{-\bar{H}_{\rm d}}=0$. However, evaluating each term in the equation is complicated.
@@ 4.4 Improving constraint fulfillment
and modified control law $u_{k}=Ew^{\rm proj}_{k}$ can be applied to ensure exact constraint fulfillment. The above is the explicit solution of $\min_{w^{\rm proj}}1/2{\left\|{w^{\rm proj}-w}\right\|}^{2}$ s.t. $h(w;x)=Cw+Dx=0$. Note that $K,L$ can be computed and stored in advance to decrease the computational burden.
@@ 5.1 Control performance
We configure a sampled-data feedback system (Figure 1(b)) that incorporates the plant described earlier and a discrete-time PDG controller $\mathcal{C}_{\rm d}$ with various parameter settings. The objective is to examine the influence of the introduced gain $\zeta$ and the sampling period $\Delta t$ on control performance. Controller parameters are set as $(\alpha,\beta)=(0.15,0.1)$, $\gamma_{k}$ is determined by Algorithm 1, and equality constraint projection in Section 4.4 is employed. In Problem 1, the control model is derived with the time-step $\Delta\tau=0.1~{}{\rm s}$. A target $x^{r}=[200/3;5]$ for the state $x$ and an upper bound $\overline{u}=160$ for the input $u$ are considered, with the objective function $f(w)={\left\|{w}\right\|}_{P}^{2}$, where $P={\rm blockdiag}(I_{2N},I_{N}/10)$ and the horizon length $N=30$. The objective function is defined with the error system configured such that the state and input are zero at the steady state of $x=x^{r},u=u^{r}\approx 133.4$. In other words, when the optimal solution is $w=0$, the system is in equilibrium at $x=x^{r},u=u^{r}$.
@@
$\delta^{*}_{\rm c}$ $0.2345$
$\lambda_{\max}^{*}(H_{\rm c})$ $-0.1055$
$\delta^{*}_{\rm d}$ $0.3447$ $0.3480$ $0.3988$ $9.7102$ $0.3981$
$\lambda_{\max}^{*}(\bar{H}_{\rm d})$ $-0.0600$ $-0.0583$ $-0.0345$ $2.0766$ $-0.0345$
@@
In the original iMPC theory, the stability of the continuous-time feedback system $(\mathcal{P}_{\rm c},\mathcal{C}_{\rm c})$ is guaranteed if $\lambda_{\max}^{*}(H_{\rm c})$ is negative. Consequently, the system is deemed stable in all cases due to $\lambda_{\max}^{*}(H_{\rm c})=-0.1055<0$, which is independent of $\zeta$ and $\Delta t$. However, as the simulations are conducted through sampled-data implementation, actual stability is compromised. In Case 4, this resulted in an infinite loop in Algorithm 1, although even one additional iteration did not occur in other cases. Note that the result for Case 4 diverged when Algorithm 1 was eliminated and $\gamma_{k}$ was fixed to 1.
By contrast, the proposed method allows for the evaluation of the impact of $\zeta,\Delta t$ on the stability of the sampled-data system through the negative definiteness of $\bar{H}_{\rm d}$. In fact, $\lambda_{\max}^{*}(\bar{H}_{\rm d})$ is positive in Case 4, indicating potential instability. Consequently, in Case 5, we designed $\lambda_{\max}^{*}(\bar{H}_{\rm d})$ to be negative by reducing $\Delta t$ with the same $\zeta$ as Case 4. Notably, $\delta^{*}_{\rm d}$ and $\lambda_{\max}^{*}(\bar{H}_{\rm d})$ in Case 5 coincide with those in Case 3. This is because $\zeta\Delta t$ exhibits the same value in both cases (see Remark 19). The control results of all cases confirmed that stability is ensured and reference tracking performance is improved by increasing the gain $\zeta$. Although it took longer for the constraint to be met in Case 3 (Figure 4), the controller was able to rapidly achieve the constraint by increasing $\zeta$ as seen in Case 5 (Figure 5). This behavior is attributed to the equilibrium condition of the controller satisfying the inequality constraint (refer to Theorem 20).
@@ 5.2 Computational efficiency
iMPC 0.574 $\leftarrow$ $\leftarrow$ 1 1 0.386 $\leftarrow$ $\leftarrow$ 1 1 0.666 $\leftarrow$ $\leftarrow$ 1 1
iMPCproj 1.289 $\leftarrow$ $\leftarrow$ 1 1 0.547 $\leftarrow$ $\leftarrow$ 1 1 1.173 $\leftarrow$ $\leftarrow$ 1 1
C/GMRES 1 7.906 $\leftarrow$ $\leftarrow$ 1 1 2.458 $\leftarrow$ $\leftarrow$ 1 1 24.24 $\leftarrow$ $\leftarrow$ 1 1
C/GMRES 2 9.149 $\leftarrow$ 1.243 2 2 3.232 $\leftarrow$ 0.774 2 2 32.14 $\leftarrow$ 7.906 2 2
Explicit MPC 1.249 0.211 0.043 29 4.902 123.7 11.78 0.105 1173 111.7 3741 436.5 0.235 15917 1857
QPKWIK 13.45 5.150 4.482 3 1.149 8.489 2.884 2.122 4 1.359 44.89 11.61 6.413 7 1.810
OSQP(warm) 110.0 5.035 4.782 23 1.053 50.42 3.882 3.878 13 1.001 142.9 5.961 5.955 24 1.001
FBRS(warm) 43.12 7.187 7.183 6 1.000 42.87 4.763 4.761 9 1.000 119.8 13.31 13.81 9 1.000
The results are summarized in Table 4. The table shows the average running time for one solver iteration (Time iter)*5*5*5The iterations for iMPC and iMPCproj represent the while loop shown in Algorithm 1, and for the Explicit MPC mean the iterations for identifying the piecewise linear region., the average time for one sampling period (Time mean), and the estimated maximum time for one sampling period (Time max), as well as the average (Iter mean) and maximum (Iter max) of the number of solver iterations in one sampling period. To eliminate the influence of unpredictable OS processes and variable CPU clock frequency, Time max is estimated by (Time mean)$\times$(Iter max)/(Iter mean).

### v3

@@ 3.1 Assumptions
For $S^{\mathcal{P}}(x):=\frac{1}{2}x^{\top}x$, there exists $Q=Q^{\top}\prec 0,\,S,\,R=R^{\top}$ such that
@@
This assumption can be satisfied with matrices such as
$H^{\mathcal{P}}_{\rm c}$ $=\begin{bmatrix}{\rm Sym}(A_{\rm c})&B_{\rm d}/2\\ B_{\rm d}^{\top}/2&O\end{bmatrix}.$
if ${\rm Sym}(A_{\rm c})\prec 0$. Even if not, you can design a pre-controller $u=Kx+v$, with new input $v$ and gain $K$, for any controllable linear systems to satisfy ${\rm Sym}(A_{\rm c}+B_{\rm c}K)\prec 0$ and just replace $A_{\rm c}$ and $u$ by $A_{\rm c}+B_{\rm c}K$ and $v$.
@@ 3.2 Instant MPC
holds for any $w,x$*1*1*1If $f(w)=\frac{1}{2}{\left\|{w}\right\|}^{2}_{P}\ (P=P^{\top},\sigma I\preceq P)$, then $\sigma I$ can be replaced as $P$ hereafter to reduce conservatism..
@@ 4.2 Assumptions
There exists matrix $H^{\mathcal{P}}_{\rm d}$, with negative-definite upper-left $n\times n$ block, and positive semi-definite matrix $P^{\mathcal{P}}_{\rm d}$ such that
@@
Assumption 11 is an extension of Assumption 4. If ${\rm Sym}(A_{\rm c})\prec 0$, this is satisfied by, e.g.,
$H^{\mathcal{P}}_{\rm d}$ $=\frac{1}{\Delta t}\begin{bmatrix}{\rm Sym}(A_{\rm d})-I&B_{\rm d}/2\\ B_{\rm d}^{\top}/2&O\end{bmatrix},$
$P^{\mathcal{P}}_{\rm d}$ $=\frac{1}{2(\Delta t)^{2}}\begin{bmatrix}A_{\rm d}-I\\ B_{\rm d}^{\top}\end{bmatrix}\begin{bmatrix}A_{\rm d}-I&B_{\rm d}\end{bmatrix}.$
@@
In addition to Assumption 8, $f$ is $\rho$–smooth convex*2*2*2A function $f(x)$ is called $\rho$–smooth convex function if it satisfies ${\left\|{\nabla{f(x)}-\nabla{f(y)}}\right\|}\leq\rho{\left\|{x-y}\right\|}$ for all $x,y$. and $\mu_{0}\geq 0$.
@@ 4.3 Controller Design
If $z_{k}\neq 0$, small $\bar{\gamma}_{k}$ satisfies (a) because $-{\left\|{z_{k}}\right\|}^{2}_{-\bar{H}_{\rm d}}<0$. If $z_{k}=0$ and $\mu_{k}\neq 0$, $\Delta V_{k}\leq\zeta\Delta t{\left(a(\mu_{k},0)\gamma_{k}+b(\mu_{k},0)\right)}\gamma_{k}$ holds from (11) and small $\bar{\gamma}_{k}$ satisfies (a) because $b(\mu_{k},0)<0$. If $z_{k}=0$ and $\mu_{k}=0$, this means $(w_{k},\mu_{k},\lambda_{k},x_{k})=(0,0,0,0)$ and (a) is trivial because $\Delta V_{k}=0\ \forall\ \gamma_{k}\in(0,1]$. To summarize what is required to guarantee stability, we need to find $\delta>0$ that satisfies (10) and determine $\gamma_{k}$ satisfying (a) at each time step. The existence of $\delta$ can be confirmed by solving an LMI, e.g., $\min_{\{\delta,{\varepsilon}\}}{\varepsilon}\ \text{s.t. }{\varepsilon}I\succeq\bar{H}_{\rm d}$. We can see from (11) that $\gamma_{k}$ can be determined by $\gamma_{k}=\min(1,c\bar{\gamma}_{k}),\ c\in(0,1)$ with deriving $\bar{\gamma}_{k}$ as the positive solution of an equation*3*3*3An example of functions $a,b$ is shown in the proof of Lemma 17 (see Appendix). Replacing $-{\left\|{z_{k}}\right\|}^{2}_{-\bar{H}_{\rm d}}$ as ${\varepsilon}{\left\|{z_{k}}\right\|}^{2}$ using the LMI solution ${\varepsilon}$ is possible to avoid a large matrix computation. $a(\mu_{k},z_{k})\gamma_{k}^{2}+b(\mu_{k},z_{k})\gamma_{k}-{\left\|{z_{k}}\right\|}^{2}_{-\bar{H}_{\rm d}}=0$. However, evaluating each term in the equation is complicated.
@@ 4.4 Improving constraint fulfillment
and modified control law $u_{k}=Ew^{\rm proj}_{k}$ can be applied to ensure exact constraint fulfillment. The above is the explicit solution of $\min_{w^{\rm proj}}1/2{\left\|{w^{\rm proj}-w}\right\|}^{2}$ s.t. $h(w^{\rm proj};x)=Cw^{\rm proj}+Dx=0$. Note that $K,L$ can be computed and stored in advance to decrease the computational burden.
@@ 5.1 Control performance
We configure a sampled-data feedback system (Figure 1(b)) that incorporates the plant described earlier and a discrete-time PDG controller $\mathcal{C}_{\rm d}$ with various parameter settings. The objective is to examine the influence of the introduced gain $\zeta$ and the sampling period $\Delta t$ on control performance. Controller parameters are set as $(\alpha,\beta)=(0.2,0.1)$, $\gamma_{k}$ is determined by Algorithm 1, and equality constraint projection in Section 4.4 is employed. In Problem 1, the control model is derived with the time-step $\Delta\tau=0.1~{}{\rm s}$. A target $x^{r}=[200/3;5]$ for the state $x$ and an upper bound $\overline{u}=160$ for the input $u$ are considered, with the objective function $f(w)={\left\|{w}\right\|}_{P}^{2}$, where $P={\rm blockdiag}(I_{2N},I_{N}/10)$ and the horizon length $N=30$. The objective function is defined with the error system configured such that the state and input are zero at the steady state of $x=x^{r},u=u^{r}\approx 133.4$. In other words, when the optimal solution is $w=0$, the system is in equilibrium at $x=x^{r},u=u^{r}$.
@@
$\delta^{*}_{\rm c}$ $0.2850$
$\lambda_{\max}^{*}(H_{\rm c})$ $-0.0870$
$\delta^{*}_{\rm d}$ $0.3662$ $0.3713$ $0.4586$ $14.7560$ $0.4578$
$\lambda_{\max}^{*}(\bar{H}_{\rm d})$ $-0.0537$ $-0.0512$ $-0.0116$ $3.2423$ $-0.0116$
@@
In the original iMPC theory, the stability of the continuous-time feedback system $(\mathcal{P}_{\rm c},\mathcal{C}_{\rm c})$ is guaranteed if $\lambda_{\max}^{*}(H_{\rm c})$ is negative. Consequently, the system is deemed stable in all cases due to $\lambda_{\max}^{*}(H_{\rm c})=-0.0870<0$, which is independent of $\zeta$ and $\Delta t$. However, as the simulations are conducted through sampled-data implementation, actual stability is compromised. In Case 4, this resulted in an infinite loop in Algorithm 1, although even one additional iteration did not occur in other cases. Note that the result for Case 4 diverged when Algorithm 1 was eliminated and $\gamma_{k}$ was fixed to 1.
By contrast, the proposed method allows for the evaluation of the impact of $\zeta,\Delta t$ on the stability of the sampled-data system through the negative definiteness of $\bar{H}_{\rm d}$. In fact, $\lambda_{\max}^{*}(\bar{H}_{\rm d})$ is positive in Case 4, indicating potential instability. Consequently, in Case 5, we designed $\lambda_{\max}^{*}(\bar{H}_{\rm d})$ to be negative by reducing $\Delta t$ with the same $\zeta$ as Case 4. Notably, $\delta^{*}_{\rm d}$ and $\lambda_{\max}^{*}(\bar{H}_{\rm d})$ in Case 5 are almost the same as in Case 3. This is because $\zeta\Delta t$ exhibits the same value in both cases (see Remark 19). The control results of all cases confirmed that stability is ensured and reference tracking performance is improved by increasing the gain $\zeta$. Although it took longer for the constraint to be met in Case 3 (Figure 4), the controller was able to rapidly achieve the constraint by increasing $\zeta$ as seen in Case 5 (Figure 5). This behavior is attributed to the equilibrium condition of the controller satisfying the inequality constraint (refer to Theorem 20).
@@ 5.2 Computational efficiency
iMPC 0.507 $\leftarrow$ $\leftarrow$ 1 1 0.386 $\leftarrow$ $\leftarrow$ 1 1 0.666 $\leftarrow$ $\leftarrow$ 1 1
iMPCproj 1.252 $\leftarrow$ $\leftarrow$ 1 1 0.547 $\leftarrow$ $\leftarrow$ 1 1 1.173 $\leftarrow$ $\leftarrow$ 1 1
C/GMRES 1 4.834 $\leftarrow$ $\leftarrow$ 1 1 2.458 $\leftarrow$ $\leftarrow$ 1 1 24.24 $\leftarrow$ $\leftarrow$ 1 1
C/GMRES 2 6.110 $\leftarrow$ 1.275 2 2 3.232 $\leftarrow$ 0.774 2 2 32.14 $\leftarrow$ 7.906 2 2
Explicit MPC 1.243 0.445 0.056 22 7.880 123.7 11.78 0.105 1173 111.7 3741 436.5 0.235 15917 1857
QPKWIK 19.70 11.67 6.565 3 1.778 8.489 2.884 2.122 4 1.359 44.89 11.61 6.413 7 1.810
OSQP(warm) 196.9 5.845 5.625 35 1.039 50.42 3.882 3.878 13 1.001 142.9 5.961 5.955 24 1.001
FBRS(warm) 66.08 9.506 9.440 7 1.007 42.87 4.763 4.761 9 1.000 119.8 13.31 13.81 9 1.000
The results are summarized in Table 4. The table shows the average running time for one solver iteration (Time iter)*4*4*4The iterations for iMPC and iMPCproj represent the while loop shown in Algorithm 1, and for the Explicit MPC mean the iterations for identifying the piecewise linear region., the average time for one sampling period (Time mean), and the estimated maximum time for one sampling period (Time max), as well as the average (Iter mean) and maximum (Iter max) of the number of solver iterations in one sampling period. To eliminate the influence of unpredictable OS processes and variable CPU clock frequency, Time max is estimated by (Time mean)$\times$(Iter max)/(Iter mean).